# Measuring Bias or Measuring the Task: Understanding the Brittle Nature of LLM Gender Biases

### v1

@@ 7 Conclusion
Large Language Models are becoming deeply integrated into social and communicative infrastructures, heightening the importance of robust, ongoing audits for harmful biases. In this work, we explore a potentially growing challenge: as these models have increasingly been exposed to past fairness evaluation and intervention data, could they show more desirable behavior when prompts look like typical gender bias evaluation formats? Our analysis provides initial evidence for this claim by finding that across models, gender-neutral pronoun use increases when we make testing- and gender-focused prompt content salient. This raises the question whether we may need to become increasingly inventive to hide our evaluative intentions when we don’t want to trigger a model’s ideal “testing mode” persona.
@@ Acknowledgments
We would like to thank the members of UCLA’s Coalas Lab for their feedback on the project. Specifically, Rohan Jain whose experiments were instrumental in inspiring the initial project idea. Further gratitude goes to our reviewers, who all provided thoughtful insights that significantly enriched the paper. We also thank Rachit Dubey and Lin L. Lin for their comments. Finally, we’re grateful that this work was financially supported by a Google Research Award.

### v3

@@ 7 Conclusion
Large Language Models are becoming deeply integrated into social and communicative infrastructures, heightening the importance of robust, ongoing audits for harmful biases. In this work, we explore a potentially growing challenge: as these models have increasingly been exposed to past fairness evaluation and intervention data, could they show more desirable behavior when prompts look like typical gender bias evaluation formats? Our analysis provides initial evidence for this claim by finding that across models, gender-neutral pronoun use increases when we make testing- and gender-focused prompt content salient. This raises the question whether we may need to become increasingly inventive to hide our evaluative intentions when we don’t want to trigger a model’s “testing mode” behavior with limited generalizability.
@@ Acknowledgments
We would like to thank the members of UCLA’s Coalas Lab for their feedback on the project. Specifically, Rohan Jain whose experiments were instrumental in inspiring the initial project idea. Further gratitude goes to our reviewers, who all provided thoughtful insights that significantly enriched the paper. We also thank Chris Potts, Rachit Dubey and Lin L. Lin for their helpful comments and guidance. Finally, we’re grateful that this work was financially supported by a Google Research Award.